A student takes a multiple-choice quiz with four questions, each of which has four different answer choices (A, B, C, and D). Assume each question has one correct answer.

Find the probability that the first two questions are answered correctly.

Answer by stanbon(75887)

A student takes a multiple-choice quiz with four questions, each of which has four different answer choices (A, B, C, and D). Assume each question has one correct answer.
Find the probability that the first two questions are answered correctly.
---
P(1st correct) = 1/4
P(2nd correct) = 1/4
----
P(1st two correct) = (1/4)^2 = 1/16
====================================
Cheers,
Stan H.